OpenStudy (anonymous):

A cat chases a mouse across a 1.5 m high table. The mouse steps out of the way, and the cat slides off the table and strikes the floor 2.1 m from the edge of the table. What was the cat’s speed when it slid off the table? The acceleration of gravity is 9.81 m/s2 . Answer in units of m/s

OpenStudy (anonymous):

I misread the question. Here's the answer Find the time to hit the floor from s = (0.5)(9.81)(t^2) =1.5m t = sqrt[(1.5)/(0.5)(9.8)] = 0.55 s constant horizontal velocity u = distance/time = 2.1 m / 0.55 s = 3.8 m/s